# Multi-intention Inverse Q-learning for Interpretable Behavior Representation

### v3

@@ 5 Experiments and discussion
We evaluate the performance of HIQL on a simulated gridworld environment and a real mice behavior dataset obtained from a $127$-node labyrinth navigation task [RZPM21], and compare to DIRL [AJP22]. We then show the potential of performing model-free learning and detecting exploration behavior with HIQL on a mice reversal-learning task [DLCSS+23].
@@ 5.1 Gridworld benchmark
We compared between the performance of HIAVI (the model-based variant of HIQL) and DIRL with $1$ or $2$ intention(s). Note that for the single intention case, HIAVI falls back to normal IAVI and DIRL falls back to maximum causal entropy IRL [ZBD10]. Trying to enable DIRL to better capture such step-like intention transition dynamics, we varied the hyperparameter that controls the smoothness of the time-varying reward in DIRL — the variance $\sigma$ for the Gaussian of the random walk prior, from $0.01$ to $10$. A larger $\sigma$ value corresponds to a less smooth intention transition dynamics in DIRL. We list the detailed information about model training in Appendix C.1.
@@
Figure 3: Results for the gridworld benchmark. (a) Comparison of HIAVI and DIRL on datasets with different number of expert demonstrations, represented as log-likelihood on the test dataset. (b) Predicted intention dynamics from HIAVI and DIRL, represented as the posterior probability of the ‘abandon’ intention and averaged across all trajectories. (c) Visualization of the ground truth and learnt state-value functions (top), and the corresponding EVDs (bottom, mean $\pm$ standard error) from HIAVI and DIRL.
@@
In general, HIAVI outperformed DIRL in predicting the expert behavior as indicated by the log-likelihood on the test dataset (Figure 3(a)). Notice that HIAVI already had a better performance compared to DIRL even in the collapsed single intention model. This difference reproduced the empirical results in [KHWB20], and indicates that IAVI serves as a better inner-loop (IRL problem) solver than maximum entropy IRL. Comparing the single- and multi-intention variants, both HIAVI and DIRL had an increase in prediction performance. Particularly, the HIAVI almost achieved the expert level as the number of demonstrations increased, while the performance improvement for DIRL was only minor. Besides, different $\sigma$ values of DIRL seem not to have a significant influence on the DIRL performance when there is enough expert demonstrations in the dataset.
@@
Next we compared HIAVI and DIRL abilities in capturing the expert’s intention transition dynamics (Figure 3(b)) and recovering the expert policy (Figure 3(c)). As a measure of performance in policy recovery, we used the expected value difference (EVD) metric [LPK11]. EVD is defined as the mean square error between the state-value under the true reward function for the expert policy and the state-value under the true reward for the optimal Boltzmann policy w.r.t. the learnt reward. It provides an estimation of the sub-optimality of the learnt policy under the true reward function. For HIAVI and DIRL with $2$ intentions, the inferred intentions were assigned to the best-fit ground truth intentions. Since the single-intention variants assume all trajectories were demonstrated under one intention, the EVD was analyzed twice on the ground truth reward for different intentions with the same learnt Boltzmann policy. Aligning with the results indicated by Figure 3(a), HIAVI had smaller EVD values than DIRL even under single-intention set up (Figure 3(c)). Then by introducing the second intention, the best performance in recovering expert policy under both intentions was achieved by HIAVI ($5.58\pm 0.47$ for $\pi^{\mathrm{goal}}$ and $6.39\pm 1.80$ for $\pi^{\mathrm{abandon}}$). While by selecting an appropriate $\sigma$ value, DIRL could reconstruct the ‘goal’ policy at very high level (with a $2.97\pm 0.67$ EVD when $\sigma=0.1$), it struggled to learn the policy under ‘abandon’ intention. As a result, the HIAVI predicted posterior probability of the expert remaining in the ‘abandon’ intention at different time steps in a single episode matched exactly with the ground truth (Figure 3(b)), whereas DIRL seemed to overestimate this probability at the beginning (where the expert reached the barrier for the first time), and underestimate it after the $8$th time step.
@@ 5.2 Real-world mice navigation benchmark
The expert demonstrations in this real-world dataset were collected from the $127$-node labyrinth navigation task. As shown in Figure 4, each black dot represents one state in the environment and the subjects can select from $4$ actions: left, right, reverse, and stay at each state. One of the $127$ states is occupied with water resource (the blue square), and the optimal path from the labyrinth entrance to the water port is depicted as blue line. Two cohorts of 10 mice moved freely in dark through the labyrinth over the course of $7$ hours, where one cohort of animals was under water restriction while the other was not. Difference in water restriction condition resulted in different animal behavior between these two cohorts in the environment. For model evaluation, HIAVI and DIRL with varied intentions from $1$ to $4$ were trained independently on the water-restricted and water-unrestricted expert demonstration datasets. We list the detailed information about trajectory preprocessing and model training in Appendix C.2.
@@
Figure 5: Results for the navigation benchmark of the water-restricted cohort. (a) Comparison of HIAVI, DIRL, and a random policy, represented as log-likelihood on the test dataset. (b) BIC as a function of the number of intentions in HIAVI. (c) Learnt policy (red arrows and crosses) in the environment and corresponding state occupancy (grey colormap) under different intentions. (d) Predicted intention dynamics from HIAVI, averaged across all trajectories. Solid and shaded curves denote the mean and standard error. (e) Inferred intention transition matrix from HIAVI.
@@
We first compared between HIAVI and DIRL on the trajectories collected from the group of water restricted animals. First of all, as the number of intentions increased, the log-likelihood value for HIAVI predictions increased significantly, whereas the improvement of DIRL performance was rather minor (Figure 5(a)). This led to a distinguishable outperformance of HIAVI compared to DIRL in this benchmark. To analyze the recovered expert policies by HIAVI, we selected the number of intentions $K=2$ according to the Bayesian information criterion (BIC) (Figure 5(b)). The learnt mice policy under intention $1$ (‘tired’) displays a preference of moving out from the water port towards the maze entrance and stay, while the policy under intention $2$ (‘thirsty’) guides the mice directly to the water port along the optimal track. Correspondingly, under the ‘tired’ intention, the highest state occupancy was observed at the entrance of the environment, while under ‘thirsty’, it was observed at the water port (Figure 5(c)). To visualize the predicted intention transition dynamics, we computed the posterior probability over mice’s intentions across all trajectories. The recovered temporal intention dynamics shows a high probability of the ‘thirsty’ intention at the beginning but later on drops gradually, as the ‘tired’ intention gradually becomes dominant (Figure 5(d)). Besides, the intention transition matrix learnt by HIAVI (Figure 5(e)) suggests that while both the ‘thirsty’ and ‘tired’ intentions are stable, the probability that the animals switch from ‘thirsty’ to ‘tired’ is relatively larger than the other way around. These observations indicate the following interpretations on the mice behavior within this cohort, which also align well with our intuition: The water-restricted mice are eager to find the water port as the trial starts, and then gradually return back to the labyrinth entrance as they obtain enough water. Once they have decided to return, the probability to go back to search for the water port would be quite low.
@@
Figure 6: Results for the navigation benchmark of the water-unrestricted cohort. (a) Comparison of HIAVI, DIRL, and a random policy, represented as log-likelihood on the test dataset. (b) BIC as a function of the number of intentions in HIAVI. (c) Learnt policy (red arrows and crosses) in the environment and corresponding state occupancy (grey colormap) under different intentions. (d) Predicted intention dynamics from HIAVI, averaged across all trajectories. Solid and shaded curves denote the mean and standard error. (e) Inferred intention transition matrix from HIAVI.
@@
When comparing the dataset collected from the water-unrestricted animals, HIAVI still had a larger log-likelihood in predicting the mice behavior than DIRL (Figure 6(a)). Interestingly, for this dataset, increasing the number of intentions seemed not to improve the performance of DIRL at all, which might be a result of a more random behavior of the animals due to the lack of motivation to find water. We then selected HIAVI with $K=2$ according to BIC (Figure 6(b)) to analyze the recovered policy and intention dynamics. For this cohort of animals, the inferred policy under two intentions exhibits ‘exploring’ and ‘tired’ behavior. The policy under ‘exploring’ tends to encourage the animal lingering in the labyrinth, whereas the policy under ‘tired’ intention steers the animal back to the maze entrance (Figure 6(c)). Correspondingly, the posterior probability of ‘exploring’ initially dominates at the beginning of the trial but is generally surpassed by the ‘tired’ intention over time (Figure 6(d)). The inferred intention transition matrix (Figure 6(e)) suggests that the ‘exploring’ intention is unstable, meaning that it has a larger probability of switching to ‘tired’ intention compared to staying at ‘exploring’. In contrast, after switching to the ‘tired’ intention, animals tend to keep this intention until the end of trial. These observations are again consistent with our intuition. Since the water-seeking motivation for this group of water-unrestricted mice is very low, they would prefer to explore the labyrinth, rather than search for the water port.
@@ 5.3 Application to mice reversal-learning behavior
Figure 7: Results for the mice reversal-learning dataset. (a) Log-likelihood (mean $\pm$ standard error, $5$-fold cross-validation) as a function of history length of single intention HIQL. (b) Change in test set log-likelihood as a function of the number of intentions in HIQL with $\ell_{h}=3$, relative to the fQ-learning model (labeled ‘F.’). Each trace represents a single mouse, averaged over cross-validation. Solid black indicates the mean across animals, and the dashed curve indicates the example mouse. (c) BIC as a function of the number of intention in HIQL with $\ell_{h}=3$. (d) Learnt mice policy represented with the probability of switch, win-stay, and lose-switch. Each grey curve denotes one mouse. (e) Average task performance and the predicted intention dynamics. Solid and shaded curves denote the mean and standard error. (f) Overall task performance (gray) and the performance under different intentions, mean $\pm$ standard error. (g) Relationship between the probability of the ‘exploitation’ intention, $5$ trials before block switch and the probability of the ‘exploration’ intention, $5$ trials after block switch, mean $\pm$ standard error.
@@
Finally, we apply the HIQL algorithm to behavioral data recorded from a group of mice engaged in a dynamic two-armed bandit reversal-learning task from [DLCSS+23]. At the beginning of the task, water-restricted mice may choose from two available spouts, left (L) and right (R), with random one of them assigned water as extrinsic reward. After reaching an online performance of $75\%$ correct in a $15$-trials sliding average window and a minimum $20$-trials block, the rewarded spout is changed. To formulate the MDP, we define the action space as: $\mathcal{A}=\{\textit{left},\textit{right}\}$. Every state $s\in\mathcal{S}$ is defined with a set of truncated history information: $s_{t}=\{\varphi_{t-1},\ldots,\varphi_{t-\ell_{h}};a_{t-1},\ldots,a_{t-\ell_{h}}\}$, where the positive integer $\ell_{h}$ denotes the history length, $a\in\mathcal{A}$ denotes history action, and $\varphi\in\{\textit{correct},\textit{error}\}$ represents trial feedback, i.e. the extrinsic reward. Such MDP formulation allows us to avoid explicitly establishing a partially observable MDP. Different from the first two experiments, we applied the model-free variant of HIQL in this dynamic reversal-learning task since the environment model is unknown.
@@
We begin from selecting the hyperparameter $\ell_{h}$ using single intention HIQL (equivalent to vanilla IQL). We compared the log-likelihood on training and test sets of multiple IQL fitting with different $\ell_{h}$ (Figure 7(a)). The log-likelihood on test sets shows a bell-shaped curve as $\ell_{h}$ increases, indicating an overfit on the training set when $\ell_{h}>3$. Note that there is an abnormal drop on training set log-likelihood at large $\ell_{h}$s. This can be explained with the insufficient sampling given the fixed set of expert demonstrations, since the size of the state space grows exponentially as the history length $\ell_{h}$ increases. The best test log-likelihood was achieved by $\ell_{h}=3$, which was selected for subsequent steps. Next, to determine the number of intentions $K$ under which mice demonstrated the trajectories, we fit multiple HIQL with varying number of intentions. In this step, we additionally applied a forgetting Q-learning (fQ-learning) model [BNLS22], which has been widely recognized as a prominent forward behavioral model for the reversal-learning task, serving as a baseline for comparative analysis. We found that the multiple intention HIQL fitting substantially outperformed the single intention models (Figure 7(b)). The BIC w.r.t. different $K$ indicates that both $K=2$ and $K=3$ are reasonable values (Figure 7(c)). Taking biological interpretability into account, we selected $K=3$ for the following discussion. We list additional details about the above procedure of fitting HIQL to this mice reversal-learning dataset in Appendix C.3.
@@
The inferred mice policies from HIQL represent different strategies in the task under three intentions (Figure 7(d)). The policy under intention $1$ (‘exploitation’) displays a strong preference of a ‘win-stay’ and ‘lose-switch’ strategy, which is the optimal policy in this deterministic reward bandit task. On the other hand, under intention $2$ (‘disengaged’), the policy exhibits a preference for exploitation when the previous trial was successful, but following error trials, it employs a random action selection, indicated by the ca. $0.5$ probability of executing a ‘lose-switch’. Lastly, under intention $3$ (‘exploration’), the subject consistently favors selecting the option opposite to the one chosen in the preceding trial, irrespective of whether they had won or lost in that particular instance. As shown in an example session (Figure 8(a)), the ‘exploration’ intention predominantly occurs at the onset of a block and only lasts for a few trials, whereas the other two intentions may persists over multiple consecutive trials. These results suggests that compared to ‘exploitation’ and ‘disengaged’, the ‘exploration’ intention is not stable, which aligns well with the learned intention transition matrix (Figure 8(b)). Additionally, it becomes evident that error trials tend to coincide with the trials where the posterior probability of ‘disengaged’ and ‘exploration’ intentions are dominant. The intention dynamics averaged across all blocks closely resembles those observed in the example session (Figures 7(e)). As each block begins with the animal’s performance at a relatively low level, there is a decline in the posterior probability associated with the ‘exploitation’ intention, accompanied by an increase in the probabilities of the other two intentions associated with suboptimal exploratory strategies. Then as the subjects’ performance steadily improves, the ‘exploitation’ intention progressively reasserts its dominance. In contrast to the cohort’s general correct rate of $0.74\pm 0.02$, the subjects performed significantly better within the ‘exploitation’ intention, achieving a correctness rate of $0.86\pm 0.01$, whereas they only attained lower correctness rates of $0.64\pm 0.03$ and $0.56\pm 0.06$ in the two other intentions (Figure 7(f)). Interestingly, it was observed that the mean posterior probability of the ‘exploration’ intention at the beginning of a new block showed a positive correlation with the average probability of the ‘exploitation’ state at the end of the preceding block (Figure 7(g)). These results confirm the difference between the two suboptimal intentions ‘disengaged’ and ‘exploration’: the first strategy explores passively upon error, while the second strategy involves a deliberate, exploration-oriented action selection when the subject is highly engaged and possess a good understanding of the environment. Although the ‘exploration’ intention is suggested to be unstable, it provides the opportunity to characterize a type of learning strategy based on error, which has been shown experimentally by Kononowicz et al. [KvWD22].
@@
In summary, by applying HIQL in a real-world mice reversal learning dataset, we demonstrate the possibility of performing model-free learning and producing interpretable behavior characteristics of animal experts via HIQL. We also detected a typical exploration strategy of mice during this value-based decision-making task via mathematical modeling, which has only be achieved in perceptual decision-making tasks [ARS+22]. Last but not the least, the model training procedure that we proposed in this experiment enables fits to individual subjects without running into the challenge of post hoc alignment of intentions from different individuals. Although the mice subjects exhibited similar behavior during our experiment, limiting the magnitude of observed differences in individual solutions, we anticipate that this method will prove beneficial in tasks where subjects exhibit greater variance in policy preferences.

### v4

@@ 5 Experiments and discussion
We evaluate the performance of HIQL on a simulated gridworld environment and a real mice behavior dataset obtained from a $127$-node labyrinth navigation task (Rosenberg et al., 2021), and compare to DIRL (Ashwood et al., 2022a). We then show the potential of performing model-free learning and detecting exploration behavior with HIQL on a mice reversal-learning task (De La Crompe et al., 2023).
@@ 5.1 Gridworld benchmark
We compared between the performance of HIAVI (the model-based variant of HIQL) and DIRL with $1$ or $2$ intention(s). Note that for the single intention case, HIAVI falls back to normal IAVI and DIRL falls back to maximum causal entropy IRL (Ziebart et al., 2010). Trying to enable DIRL to better capture such step-like intention transition dynamics, we varied the hyperparameter that controls the smoothness of the time-varying reward in DIRL — the variance $\sigma$ for the Gaussian of the random walk prior, from $0.01$ to $10$. A larger $\sigma$ value corresponds to a less smooth intention transition dynamics in DIRL. The whole expert demonstration dataset consisted of $1024$ trajectories. All evaluated algorithms were fit to multiple sub-datasets with different number of expert trajectories, and each dataset was analyzed using a $5$-fold cross-validation. We list the detailed information about model training in Appendix C.1.
@@
Figure 3: Results for the gridworld benchmark. (a) Comparison of HIAVI and DIRL on datasets with different number of expert trajectories, represented as log-likelihood on the test dataset (mean $\pm$ standard error, 5-fold cross-validation). (b) Predicted intention dynamics from HIAVI and DIRL using the outputs from the best cross-validation fold, represented as the posterior probability of the ‘abandon’ intention and averaged across all trajectories (mean $\pm$ standard error across $1024$ trajectories). (c) Visualization of the ground truth and learnt state-value functions from the best cross-validation fold (top), and the corresponding EVDs (bottom, mean $\pm$ standard error, 5-fold cross-validation) from HIAVI and DIRL.
@@
In general, HIAVI outperformed DIRL in predicting the expert behavior as indicated by the log-likelihood on the test dataset (Figure 3(a)). Notice that HIAVI already had a better performance compared to DIRL even in the collapsed single intention model, indicating that IAVI serves as a better inner-loop (IRL problem) solver than maximum entropy IRL. This is probably because IAVI learns the unknown expert reward in closed-form by solving a system of linear equations instead of performing parameter estimation as maximum entropy IRL does, which introduces approximate error (Kalweit et al., 2020). Comparing the single- and multi-intention variants, both HIAVI and DIRL had an increase in prediction performance. Particularly, the HIAVI almost achieved the expert level as the number of trajectories increased, while the performance improvement for DIRL was only minor. Besides, different $\sigma$ values of DIRL seem not to have a significant influence on the DIRL performance when there is enough expert trajectories in the dataset.
@@
Next we compared HIAVI and DIRL abilities in capturing the expert’s intention transition dynamics (Figure 3(b)) and recovering the expert policy (Figure 3(c)). The analysis was performed based on the learnt policy using the whole dataset with all $1024$ trajectories (with cross-validation applied). As a measure of performance in policy recovery, we used the expected value difference (EVD) metric (Levine et al., 2011). EVD is defined as the mean square error between the state-value under the true reward function for the expert policy and the state-value under the true reward for the optimal Boltzmann policy w.r.t. the learnt reward. It provides an estimation of the sub-optimality of the learnt policy under the true reward function. For HIAVI and DIRL with $2$ intentions, the inferred intentions were assigned to the best-fit ground truth intentions, and the two learnt policy was evaluated under the corresponding true reward individually to obtain the respective state-values. Under the single-intention setup, since only one learnt policy could be provided by the algorithms, the same output learnt policy was evaluated twice repeatedly under the different true reward functions, respectively. Aligning with the results indicated by Figure 3(a), HIAVI had smaller EVD values than DIRL even under single-intention setup (Figure 3(c)). Particularly, the learnt policy from single-intention HIAVI exhibits a mixed characteristic of both ‘goal’ and ‘thirsty’ intentions but is more close to the former, whereas the learnt policy from single-intention DIRL deviates from both of the two intentions. This explains why HIAVI already had a better prediction log-likelihood than DIRL under the single-intention setup. Then by introducing the second intention, the best performance in recovering expert policy under both intentions was achieved by HIAVI ($5.58\pm 0.47$ for $\pi^{\mathrm{goal}}$ and $6.39\pm 1.80$ for $\pi^{\mathrm{abandon}}$). While by selecting an appropriate $\sigma$ value, DIRL could reconstruct the ‘goal’ policy at very high level (with a $2.97\pm 0.67$ EVD when $\sigma=0.1$), it struggled to learn the policy under ‘abandon’ intention. As a result, the HIAVI predicted posterior probability of the expert remaining in the ‘abandon’ intention at different time steps in a single episode matched exactly with the ground truth (Figure 3(b)), whereas DIRL seemed to overestimate this probability at the beginning (where the expert reached the barrier for the first time), and underestimate it after the $8$th time step.
@@ 5.2 Real-world mice navigation benchmark
The expert demonstrations in this real-world dataset were originally collected by Rosenberg et al. (2021) from a $127$-node labyrinth navigation task. As shown in Figure 4, each black dot represents one state in the environment and the subjects can select from $4$ actions: left, right, reverse, and stay at each state. One of the $127$ states is occupied with water resource (the blue square), and the optimal path from the labyrinth entrance to the water port is depicted as blue line. Two cohorts of 10 mice moved freely in dark through the labyrinth over the course of $7$ hours, where one cohort of animals was under water restriction while the other was not. Difference in water restriction condition resulted in different animal behavior between these two cohorts in the environment. For model evaluation, HIAVI and DIRL with varied intentions from $1$ to $4$ were trained independently on the water-restricted and water-unrestricted expert demonstration datasets (with $200$ and $207$ animal trajectories, respectively). $20\%$ of the trajectories from each dataset were held out as a test set. We list the detailed information about trajectory preprocessing and model training in Appendix C.2.
@@
Figure 5: Results for the navigation benchmark of the water-restricted cohort. (a) Comparison of HIAVI, DIRL, and a random policy, represented as log-likelihood on the test dataset. (b) BIC as a function of the number of intentions in HIAVI. (c) Learnt policy (red arrows and crosses) in the environment and corresponding state occupancy (grey colormap) under different intentions. (d) Predicted intention dynamics from HIAVI, averaged across all $200$ trajectories. Solid and shaded curves denote the mean and standard error across $200$ trajectories. (e) Inferred intention transition matrix from HIAVI.
@@
We first compared between HIAVI and DIRL on the trajectories collected from the group of water restricted animals. First of all, as the number of intentions increased, the log-likelihood value for HIAVI predictions increased significantly, whereas the improvement of DIRL performance was rather minor (Figure 5(a)). This led to a distinguishable outperformance of HIAVI compared to DIRL in this benchmark. To check whether the better performance of HIAVI compared to DIRL was due to a larger number of papameters, we then calculated the Bayesian information criterion (BIC) for these two approaches (Figure 5(b)). It shows that the BIC values for HIAVI are consistently lower than those for DIRL when the number of intentions exceeds one. To analyze the recovered expert policies by HIAVI, we selected the number of intentions $K=2$ where the lowest BIC value was achieved. The learnt mice policy under intention $1$ (‘tired’) displays a preference of moving out from the water port towards the maze entrance and stay, while the policy under intention $2$ (‘thirsty’) guides the mice directly to the water port along the optimal track. Correspondingly, under the ‘tired’ intention, the highest state occupancy was observed at the entrance of the environment, while under ‘thirsty’, it was observed at the water port (Figure 5(c)). To visualize the predicted intention transition dynamics, we computed the posterior probability over mice’s intentions across all trajectories. The recovered temporal intention dynamics shows a high probability of the ‘thirsty’ intention at the beginning but later on drops gradually, as the ‘tired’ intention gradually becomes dominant (Figure 5(d)). Besides, the intention transition matrix learnt by HIAVI (Figure 5(e)) suggests that while both the ‘thirsty’ and ‘tired’ intentions are stable, the probability that the animals switch from ‘thirsty’ to ‘tired’ is relatively larger than the other way around. These observations indicate the following interpretations on the mice behavior within this cohort, which also align well with our intuition: The water-restricted mice are eager to find the water port as the trial starts, and then gradually return back to the labyrinth entrance as they obtain enough water. Once they have decided to return, the probability to go back to search for the water port would be quite low.
@@
Figure 6: Results for the navigation benchmark of the water-unrestricted cohort. (a) Comparison of HIAVI, DIRL, and a random policy, represented as log-likelihood on the test dataset. (b) BIC as a function of the number of intentions in HIAVI. (c) Learnt policy (red arrows and crosses) in the environment and corresponding state occupancy (grey colormap) under different intentions. (d) Predicted intention dynamics from HIAVI, averaged across all $207$ trajectories. Solid and shaded curves denote the mean and standard error across $207$ trajectories. (e) Inferred intention transition matrix from HIAVI.
@@
When comparing the dataset collected from the water-unrestricted animals, HIAVI also had a larger test log-likelihood and a lower BIC value than DIRL (Figure 6(a) and 6(b)). Interestingly, for this dataset, increasing the number of intentions seemed not to improve the performance of DIRL at all, which might be a result of a more random behavior of the animals due to the lack of motivation to find water. We then selected HIAVI with $K=2$ according to BIC (Figure 6(b)) to analyze the recovered policy and intention dynamics. For this cohort of animals, the inferred policy under two intentions exhibits ‘exploring’ and ‘tired’ behavior. The policy under ‘exploring’ tends to encourage the animal lingering in the labyrinth, whereas the policy under ‘tired’ intention steers the animal back to the maze entrance (Figure 6(c)). Correspondingly, the posterior probability of ‘exploring’ initially dominates at the beginning of the trial but is generally surpassed by the ‘tired’ intention over time (Figure 6(d)). The inferred intention transition matrix (Figure 6(e)) suggests that the ‘exploring’ intention is unstable, meaning that it has a larger probability of switching to ‘tired’ intention compared to staying at ‘exploring’. In contrast, after switching to the ‘tired’ intention, animals tend to keep this intention until the end of trial. These observations are again consistent with our intuition. Since the water-seeking motivation for this group of water-unrestricted mice is very low, they would prefer to explore the labyrinth, rather than search for the water port.
@@ 5.3 Application to mice reversal-learning behavior
Figure 7: Results for the mice reversal-learning dataset. (a) Log-likelihood (mean $\pm$ standard error, $5$-fold cross-validation) as a function of history length of single intention HIQL. (b) Change in test set log-likelihood as a function of the number of intentions in HIQL with $\ell_{h}=3$, relative to the fQ-learning model (labeled ‘F.’). Each trace represents a single mouse, averaged over cross-validation. Solid black indicates the mean across animals, and the dashed curve indicates the example mouse. (c) BIC as a function of the number of intention in HIQL with $\ell_{h}=3$. (d) Learnt mice policy represented with the probability of switch, win-stay, and lose-switch. Each grey curve denotes one mouse. (e) Average task performance and the predicted intention dynamics. Solid and shaded curves denote the mean and standard error across $9$ animals. (f) Overall task performance (gray) and the performance under different intentions, mean $\pm$ standard error across $9$ animals. (g) Relationship between the probability of the ‘exploitation’ intention, $5$ trials before block switch and the probability of the ‘exploration’ intention, $5$ trials after block switch, mean $\pm$ standard error across $9$ animals.
@@
Finally, we apply the HIQL algorithm to behavioral data recorded from a group of mice engaged in a dynamic two-armed bandit reversal-learning task. The dataset was initially reported by De La Crompe et al. (2023). At the beginning of the task, water-restricted mice may choose from two available spouts, left (L) and right (R), with random one of them assigned water as extrinsic reward. After reaching an online performance of $75\%$ correct in a $15$-trials sliding average window and a minimum $20$-trials block, the rewarded spout is changed. To formulate the MDP, we define the action space as: $\mathcal{A}=\{\textit{left},\textit{right}\}$. Every state $s\in\mathcal{S}$ is defined with a tuple of truncated history information: $s_{t}=(\varphi_{t-1},\ldots,\varphi_{t-\ell_{h}};a_{t-1},\ldots,a_{t-\ell_{h}})$, where the positive integer $\ell_{h}$ denotes the history length, $a\in\mathcal{A}$ denotes history action, and $\varphi\in\{\textit{correct},\textit{error}\}$ represents trial feedback, i.e., the extrinsic reward. Such MDP formulation allows us to avoid explicitly establishing a partially observable MDP. Different from the first two experiments, we applied the model-free variant of HIQL in this dynamic reversal-learning task since the environment model is unknown.
@@
We begin from selecting the hyperparameter $\ell_{h}$ using single intention HIQL (equivalent to vanilla IQL). The dataset for all subsequent analysis consists of $64$ sessions (trajectories) from $9$ different animals and a $5$-fold cross-validation was applied over the trajectories. We compared the log-likelihood on training and test sets of multiple IQL fitting with different $\ell_{h}$ (Figure 7(a)). The log-likelihood on test sets shows a bell-shaped curve as $\ell_{h}$ increases, indicating an overfit on the training set when $\ell_{h}>3$. Note that there is an abnormal drop on training set log-likelihood at large $\ell_{h}$s. This can be explained with the insufficient sampling given the fixed set of expert demonstrations, since the size of the state space grows exponentially as the history length $\ell_{h}$ increases. The best test log-likelihood was achieved by $\ell_{h}=3$, which was selected for subsequent steps. Next, to determine the number of intentions $K$ under which mice demonstrated the trajectories, we fit multiple HIQL with varying number of intentions. In this step, we additionally applied a forgetting Q-learning (fQ-learning) model (Beron et al., 2022), which has been widely recognized as a prominent forward behavioral model for the reversal-learning task, serving as a baseline for comparative analysis. We found that the multiple intention HIQL fitting substantially outperformed the single intention models (Figure 7(b)). The BIC w.r.t. different $K$ indicates that both $K=2$ and $K=3$ are reasonable values (Figure 7(c)). Taking biological interpretability into account, we selected $K=3$ for the following discussion. We list additional details about the above procedure of fitting HIQL to this mice reversal-learning dataset in Appendix C.3.
@@
The inferred mice policies from HIQL represent different strategies in the task under three intentions (Figure 7(d)). The policy under intention $1$ (‘exploitation’) displays a strong preference of a ‘win-stay’ and ‘lose-switch’ strategy, which is the optimal policy in this deterministic reward bandit task, i.e., stay on the same side for the next trial if it is rewarded in the current trial, and switch to the other one otherwise. On the other hand, under intention $2$ (‘disengaged’), the policy exhibits a preference for exploitation when the previous trial was successful, but following error trials, it employs a random action selection, indicated by the ca. $0.5$ probability of executing a ‘lose-switch’. Lastly, under intention $3$ (‘exploration’), the subject consistently favors selecting the option opposite to the one chosen in the preceding trial, irrespective of whether they had won or lost in that particular instance. As shown in an example session (Figure 8(a)), the ‘exploration’ intention predominantly occurs at the onset of a block and only lasts for a few trials, whereas the other two intentions may persists over multiple consecutive trials. These results suggests that compared to ‘exploitation’ and ‘disengaged’, the ‘exploration’ intention is not stable, which aligns well with the learned intention transition matrix (Figure 8(b)). Additionally, it becomes evident that error trials tend to coincide with the trials where the posterior probability of ‘disengaged’ and ‘exploration’ intentions are dominant. The intention dynamics averaged across all blocks closely resembles those observed in the example session (Figures 7(e)). As each block begins with the animal’s performance at a relatively low level, there is a decline in the posterior probability associated with the ‘exploitation’ intention, accompanied by an increase in the probabilities of the other two intentions associated with suboptimal exploratory strategies. Then as the subjects’ performance steadily improves, the ‘exploitation’ intention progressively reasserts its dominance. In contrast to the cohort’s general correct rate of $0.74\pm 0.02$, the subjects performed significantly better within the ‘exploitation’ intention, achieving a correctness rate of $0.86\pm 0.01$, whereas they only attained lower correctness rates of $0.64\pm 0.03$ and $0.56\pm 0.06$ in the two other intentions (Figure 7(f)). Interestingly, it was observed that the mean posterior probability of the ‘exploration’ intention at the beginning of a new block showed a positive correlation with the average probability of the ‘exploitation’ state at the end of the preceding block (Figure 7(g)). These results confirm the difference between the two suboptimal intentions ‘disengaged’ and ‘exploration’: the first strategy explores passively upon error, while the second strategy involves a deliberate, exploration-oriented action selection when the subject is highly engaged and possess a good understanding of the environment. Although the ‘exploration’ intention is suggested to be unstable, it provides the opportunity to characterize a type of learning strategy based on error, which has been shown experimentally by Kononowicz et al. (2022).
@@
In summary, by applying HIQL in a real-world mice reversal learning dataset, we demonstrate the possibility of performing model-free learning and producing interpretable behavior characteristics of animal experts via HIQL. We also detected a typical exploration strategy of mice during this value-based decision-making task via mathematical modeling, which has only be achieved in perceptual decision-making tasks (Ashwood et al., 2022b). Last but not the least, the model training procedure that we proposed in this experiment enables fits to individual subjects without running into the challenge of post hoc alignment of intentions from different individuals. Although the mice subjects exhibited similar behavior during our experiment, limiting the magnitude of observed differences in individual solutions, we anticipate that this method will prove beneficial in tasks where subjects exhibit greater variance in policy preferences.